Determining missing items in return and residual income computations Data are presented in the following table of returns on分分分
Answer #1
a. Retun on Investment = 257400/990000=26%
b. $990000X 15%=$148500
c. 257400-148500= $108900
d. $64900+29500= $94400
e. 94400/590000= 16%
f. 64900/590000= 11%
g. 350000X 14%= $49000
h. 35000/350000= 10%
i. $49000-35000=$14000
j. 54600/260000= 21%
k. 260000X 13%= $33800
l. $54600-$33800= $20800